10. (a) For a distribution, the mean is 10, variance is 16, the skewness 4

sk is +1 and kurtosis 

b2

 is 4. Obtain the first four moments about the origin i.e. zero. Comment upon the 

nature of the distribution.


1
Expert's answer
2021-03-30T07:08:13-0400

Mean "= m_1\u2032 =10" , variance ="m_2 =16" , coefficient of skewness "= \gamma_1 = +1" , coefficient of kurtosis is "\beta_2 = 4."

First moment about the origin is

Mean "= m_1\u2032 = 10" .


Next, Variance "= 𝑚_2 = 16 =\dfrac{∑ x}{n}-(\dfrac{\sum x}{n})^2"


Second moment about the origin is "m_2=\dfrac{\sum x^2}{n}=m_2+\mu_1^2=Variance+(\dfrac{\sum x}{n})^2=16+10^2=116"


Coefficient of skewness is +1:

"\gamma1 = +1 =\dfrac{𝑚_3}{𝑚_2^{\frac{3}{2}}}\rightarrow 𝑚_3 = 𝑚_2^{\frac{3}{2}} = 16^{\frac{3}{2}} = 64"

    

Coefficient of kurtosis is 4:


"\beta_2 = 4\\

\dfrac{𝑚_4}{𝑚_2^2} = 4 \rightarrow 𝑚_4 = 4(16)^2 = 1024"

Third moment about the origin is


"\ud835\udc5a_3\u2032 = \ud835\udc5a_3 + 3\ud835\udc5a_2\ud835\udc5a_1\u2032 + \ud835\udc5a_1\u2032 ^3 = 64 + 3(16)(10) + 10^3 = \ud835\udfcf\ud835\udfd3\ud835\udfd2\ud835\udfd2"


Fourth moment about the origin is


"\ud835\udc5a_4\u2032 = \ud835\udc5a_4 + 4\ud835\udc5a_3\ud835\udc5a_1\u2032 + 6\ud835\udc5a_2\ud835\udc5a_2\u2032 + \ud835\udc5a_1\u2032 ^4 = 1024 + 4(64)(10) + 6(16)102 + 10^4 = \ud835\udfd0\ud835\udfd1\ud835\udfcf\ud835\udfd6\ud835\udfd2"


It is a positively skewed distribution.